# PASER [Uncaptioned image]: Post-Training Data Selection for Efficient Pruned Large Language Model Recovery

### v1

@@ Appendix B Pseudocode
  end function
Appendix C Evaluation on Mathematical Problem Solving Tasks
  • GSM8K (Cobbe et al., 2021): A dataset containing 8.5K high-quality grade school math word problems that test various mathematical problem solving capabilities, including arithmetic, algebra, and word problem solving.
  • Minerva Math (Lewkowycz et al., 2022): A comprehensive mathematical evaluation dataset covering diverse topics in mathematics ranging from arithmetic to calculus, with problems requiring multi-step reasoning.
@@ Appendix C Evaluation on Mathematical Problem Solving Tasks
Appendix D Evaluation on Code Generation Tasks
@@
  • HumanEval (Chen et al., 2021): A widely-used code generation benchmark consisting of 164 hand-crafted Python programming problems that test various programming concepts. Each problem contains a function signature with docstring and test cases, requiring models to complete the implementation. The benchmark evaluates functional correctness using the Pass@$k$ metric, which measures the percentage of problems where a correct solution appears in k samples.
  • MBPP (Austin et al., 2021): A programming benchmark containing 974 Python programming problems focused on basic programming tasks. Each problem includes a natural language description, test cases, and a reference solution, making it particularly suitable for evaluating language-to-code generation capabilities. MBPP uses the same Pass@$k$ evaluation metric as HumanEval but generally features simpler problems with a broader coverage of basic programming concepts.
@@
Appendix E Extended Experiments on Post-quantization Recovery Training
Though the method descriptions and the experiments in the main body are mainly around the LLM pruning scenario, our PASER framework can actually be extended seamlessly to other LLM compression scenario. To further demonstrate its applicability, we conduct the experiments on post-quantization recovery training and compare our PASER with corresponding instruction tuning data selection baselines. In detail, we choose two most representative methods: Round-To-Nearest (RTN) (Frantar & Alistarh, 2022; Yao et al., 2022), GPTQ (Frantar et al., 2023), to perform the LLM quantization. It should be clarified that RTN method, which rounds all weights to the nearest quantized value on exactly the same asymmetric per-row grid, is actually the fundamental technique in most works about LLM quantization (Frantar & Alistarh, 2022; Yao et al., 2022; Park et al., ). Its runtimes scales well to the models with many billion parameters due to the direct rounding. According to the results provided in Table 7 and 8, we can observe that the PASER can still effectively enhance the recovery performance and outperform the data selection baselines on averaged reasoning performance and zero-shot perplexity for both LLaMA2-7B and LLaMA2-13B models. Meanwhile, recovery data selection baselines can indeed achieve the stronger performance than full data and random baselines, which validates the necessity of conducting recovery data selection even in the LLM quantization scenario. Furthermore, comparing these results with Table 1 and 2, it can be noticed that the improvement space of PASER in Table 7 and 8 has been reduced to some extent. This is because the post-compression performance of such quantization schemes has been competitive enough, which can reflected from the w/o training row.
@@ Appendix E Extended Experiments on Post-quantization Recovery Training
Appendix F Extended Experiments on Recovery Training with Knowledge Distillation
Inspired by  (Muralidharan et al., 2024), we explore the knowledge distillation as the recovery post-training paradigm instead of the standard supervised learning with the groundtruth label. Here, we set the original model $M_{o}$ as the teacher and the pruned model $M_{p}$ as the student. The mean KL divergence (Kullback & Leibler, 1951) between the output probability distribution of $M_{o}$ and that of $M_{p}$ is taken as the objective function. Comparing the corresponding results under different pruning schemes in Table 9 with that in Table 1, we can first observe that knowledge distillation can effectively improve the recovery performance on both reasoning and language modeling tasks in most cases. In particular, the reasoning performance of PASER is improved by 0.348 points on average among such four pruning schemes. Interestingly, the knowledge distillation recovery performance of Full Data under LLM-Pruner is much better than that with standard label-supervised learning. This demonstrates that knowledge distillation is also a promising approach to avoid the misleading information from the irrelevant or conflicting samples existing in the original dataset. Because its learning process directly imitates the unpruned model behavior instead of the provided labels, thus better preserving the thinking and decision-making consistency with the original model. As a summary, distilling the knowledge of unpruned model into the pruned model can be regarded as an effective way to enhance the recovery performance, though bring more memory overhead. Furthermore, stronger layer-wise distillation can also be taken into consideration (Jiao et al., 2020).

### v2

@@ Appendix B Pseudocode
end procedure
Appendix C Theoretical Error Bound Analysis
In this section, we establish theoretical guarantees for our PASER framework by deriving error bounds that characterize how close PASER’s solution is to the optimal recovery solution.
C.1 Problem Formulation Revisited
Recall from Section 3.1 that our objective is to find a subset $S^{*}\subset D$ of instruction tuning data that minimizes the expected loss on downstream tasks:
$S^{*}=\arg\min_{S\subset D,|S|\leq B}\mathbb{E}_{(x,y)\sim\mathcal{T}}[\mathcal{L}(M_{r}(S),x,y)],$ (8)
where $M_{r}(S)$ is the recovered model after training on subset $S$, $\mathcal{T}$ is the distribution of downstream evaluation tasks, and $\mathcal{L}$ is a loss function.
C.2 Theoretical Framework
To establish theoretical guarantees, we introduce the following assumptions:
Assumption 1 (Performance Decomposition).
The performance on downstream tasks can be decomposed as a weighted sum of performances across different capability clusters:
$\mathbb{E}_{(x,y)\sim\mathcal{T}}[L(M_{r}(S),x,y)]=\sum_{k=1}^{K}w_{k}\cdot\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S),x,y)],$ (9)
where $c_{k}$ represents the $k$-th capability cluster, $w_{k}\geq 0$ is its weight in the evaluation distribution, and $\sum_{k=1}^{K}w_{k}=1$.
Assumption 2 (Capability Independence).
The recovery effect on one capability depends primarily on samples from the corresponding cluster, with limited cross-cluster effects. For distinct clusters $c_{i}$ and $c_{j}$, and subsets $S_{i}\subset c_{i}$ and $S_{j}\subset c_{j}$:
$|\mathbb{E}_{(x,y)\sim c_{i}}[L(M_{r}(S_{i}\cup S_{j}),x,y)]-\mathbb{E}_{(x,y)\sim c_{i}}[L(M_{r}(S_{i}),x,y)]|\leq\delta_{ij},$ (10)
where $\delta_{ij}\geq 0$ is a small constant representing the cross-cluster influence.
Assumption 3 (Conceptual Consistency Monotonicity).
Within a capability cluster, for conceptually consistent subsets, larger training subsets generally produce better recovery effects. Given a Concept Consistency Graph $G=(V,E)$, for any $S_{A}\subset S_{B}\subset c_{k}$ where all pairs of samples in $S_{B}$ are conceptually consistent according to $G$:
$\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{B}),x,y)]\leq\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{A}),x,y)].$ (11)
Assumption 4 (Sample Selection Submodularity).
The recovery benefit exhibits approximate submodularity within each capability cluster. For conceptually consistent subsets $S\subset T\subset c_{k}$ and an element $e\in c_{k}\setminus T$ that is conceptually consistent with all elements in $T$:
$\Delta(e|S)\geq\Delta(e|T)-\alpha_{k}\cdot|T\setminus S|,$ (12)
where $\Delta(e|S)=\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S),x,y)]-\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S\cup\{e\}),x,y)]$ represents the marginal benefit of adding element $e$ to subset $S$, and $\alpha_{k}>0$ is a cluster-specific constant.
Assumption 5 (Intra-Capability Selection Effectiveness).
For each capability cluster $c_{k}$, let $S_{k}$ be the subset of size $n_{k}$ selected by PASER and $S_{k}^{*}$ be the optimal subset of the same size. The intra-capability selection effectiveness of PASER is characterized by:
$\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}),x,y)]\leq\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}^{*}),x,y)]+(1-\gamma_{k})\cdot\Delta_{k},$ (13)
where $\gamma_{k}\in(0,1]$ and $\Delta_{k}=\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(\emptyset),x,y)]-\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}^{*}),x,y)]$ represents the maximum possible improvement, which is guaranteed by above Assumption 4.
Assumption 6 (Inter-Capability Budget Allocation Effectiveness).
Given the limited cross-cluster influence established in Assumption 2, suboptimal budget allocation across capability clusters cannot catastrophically impact overall recovery performance. Therefore, the effectiveness of PASER’s budget allocation across clusters can be bounded by:
$\sum_{k=1}^{K}w_{k}\cdot\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}^{*}),x,y)]\leq\sum_{k=1}^{K}w_{k}\cdot\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}^{*}(n_{k}^{*})),x,y)]+(1-\beta)\cdot C_{1},$ (14)
where $\beta\in(0,1]$, $S_{k}^{*}$ is the optimal subset of size $n_{k}$ from cluster $c_{k}$, $S_{k}^{*}(n_{k}^{*})$ is the optimal subset of size $n_{k}^{*}$ from cluster $c_{k}$, and $\{n_{k}^{*}\}_{k=1}^{K}$ is the optimal budget allocation across clusters under the same total budget $B$ ($\sum_{k=1}^{K}n_{k}=\sum_{k=1}^{K}n_{k}^{*}=B$). Thus, $S_{k}^{*}(n_{k}^{*})=S^{*}\cap c_{k}$.
Based on these assumptions, we establish the following theorem:
Theorem 2 (PASER Error Bound).
Let $S_{PASER}$ be the subset selected by PASER with budget $B$, $S^{*}$ be the optimal subset of the same size, and $\gamma=\min_{k}\gamma_{k}$. Then there exists a constant $C>0$ such that:
$\mathbb{E}_{(x,y)\sim\mathcal{T}}[L(M_{r}(S_{PASER}),x,y)]\leq\mathbb{E}_{(x,y)\sim\mathcal{T}}[L(M_{r}(S^{*}),x,y)]+(1-\beta\gamma)\cdot C.$ (15)
C.3 Detailed Derivation
We now provide a step-by-step derivation of the error bound, explicitly showing how each assumption is utilized.
Step 1: Decompose the performance into capability clusters using Assumption 1:
$\mathbb{E}_{(x,y)\sim\mathcal{T}}[L(M_{r}(S_{PASER}),x,y)]$ $=\sum_{k=1}^{K}w_{k}\cdot\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{PASER}),x,y)]$ (16)
$\mathbb{E}_{(x,y)\sim\mathcal{T}}[L(M_{r}(S^{*}),x,y)]$ $=\sum_{k=1}^{K}w_{k}\cdot\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S^{*}),x,y)].$ (17)
Step 2: Apply Assumption 2 (Capability Independence) to isolate the effect of samples from each cluster:
$\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{PASER}),x,y)]$ $\leq\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{PASER}\cap c_{k}),x,y)]+\sum_{j\neq k}\delta_{kj}$ (18)
$=\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}),x,y)]+\sum_{j\neq k}\delta_{kj},$ (19)
where $S_{k}=S_{PASER}\cap c_{k}$ is the subset selected by PASER from cluster $c_{k}$.
Similarly for the optimal subset:
$\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S^{*}),x,y)]\geq\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S^{*}\cap c_{k}),x,y)]-\sum_{j\neq k}\delta_{kj}.$ (20)
Step 3: Now we analyze the relationship between PASER’s selection $S_{k}$ and the optimal selection $S_{k}^{*}$ within cluster $c_{k}$.
First, we utilize Assumption 3 (Conceptual Consistency Monotonicity). PASER’s Concept Consistency Graph ensures that all selected samples in $S_{k}$ are conceptually consistent. For any subset $S_{k}^{CCG}\subset S_{k}$ that is conceptually consistent:
$\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}),x,y)]\leq\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}^{CCG}),x,y)].$ (21)
Now, we apply Assumption 4 (Sample Selection Submodularity) to analyze the greedy selection process employed by PASER within each cluster. For a conceptually consistent set, the greedy algorithm achieves an approximation ratio that depends on the degree of submodularity.
Let $S_{k}^{G}$ be the set obtained by a greedy selection process that maximizes the marginal gain at each step while ensuring conceptual consistency. From the properties of approximated submodular functions (Assumption 4) and the guarantees of the greedy algorithm for such functions:
$\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}^{G}),x,y)]\leq\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}^{*}),x,y)]+\eta_{k}\cdot\Delta_{k},$ (22)
where $\eta_{k}\in[0,1)$ is a constant that depends on the submodularity gap parameter $\alpha_{k}$ and the size of the selection $n_{k}$, and $\Delta_{k}$ is as defined in Assumption 5 (Intra-Capability Selection Effectiveness).
PASER’s efficiency-driven selection further refines this greedy selection by prioritizing samples with higher Individual Efficiency Scores. This introduces an additional approximation factor $\phi_{k}\in[0,1)$:
$\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}),x,y)]\leq\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}^{G}),x,y)]+\phi_{k}\cdot\Delta_{k}.$ (23)
Combining these equations, we get:
$\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}),x,y)]\leq\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}^{*}),x,y)]+(\eta_{k}+\phi_{k})\cdot\Delta_{k}.$ (24)
Setting $(1-\gamma_{k})=(\eta_{k}+\phi_{k})$, we arrive at the form presented in Assumption 5:
$\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}),x,y)]\leq\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}^{*}),x,y)]+(1-\gamma_{k})\cdot\Delta_{k}.$ (25)
This explicitly shows how $\gamma_{k}$ incorporates both the submodularity properties (Assumption 4) and the conceptual consistency properties (Assumption 3) into our error bound.
Step 4: Combine Steps 2 and 3:
$\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{PASER}),x,y)]$ $\leq\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}),x,y)]+\sum_{j\neq k}\delta_{kj}$ (26)
$\leq\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}^{*}),x,y)]+(1-\gamma_{k})\cdot\Delta_{k}+\sum_{j\neq k}\delta_{kj}.$ (27)
Step 5: Apply Assumption 5 (Inter-Capability Budget Allocation Optimality) to compare PASER’s allocation with the optimal allocation:
$\sum_{k=1}^{K}w_{k}\cdot\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}^{*}),x,y)]$ $\leq\sum_{k=1}^{K}w_{k}\cdot\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}^{*}(n_{k}^{*})),x,y)]+(1-\beta)\cdot C_{1},$ (28)
where $C_{1}$ is a constant related to the maximum performance difference between different budget allocations. The factor $\beta$ quantifies how well PASER’s capability degradation-aware budget allocation approximates the theoretical optimal allocation.
Step 6: Combine Steps 1-5 and let $\gamma=\min_{k}\gamma_{k}$:
$\mathbb{E}_{(x,y)\sim\mathcal{T}}[L(M_{r}(S_{PASER}),x,y)]=\sum_{k=1}^{K}w_{k}\cdot\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{PASER}),x,y)]$ (29)
$\leq\sum_{k=1}^{K}w_{k}\cdot\left(\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}^{*}),x,y)]+(1-\gamma_{k})\cdot\Delta_{k}+\sum_{j\neq k}\delta_{kj}\right)$ (30)
$\leq\sum_{k=1}^{K}w_{k}\cdot\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}^{*}),x,y)]+(1-\gamma)\cdot\sum_{k=1}^{K}w_{k}\cdot\Delta_{k}+\sum_{k=1}^{K}w_{k}\sum_{j\neq k}\delta_{kj}$ (31)
$\leq\sum_{k=1}^{K}w_{k}\cdot\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S_{k}^{*}(n_{k}^{*})),x,y)]+(1-\beta)\cdot C_{1}+(1-\gamma)\cdot C_{2}+C_{3}$ (32)
$=\sum_{k=1}^{K}w_{k}\cdot\mathbb{E}_{(x,y)\sim c_{k}}[L(M_{r}(S^{*}\cap c_{k}),x,y)]+(1-\beta)\cdot C_{1}+(1-\gamma)\cdot C_{2}+C_{3}$ (33)
$\leq\mathbb{E}_{(x,y)\sim\mathcal{T}}[L(M_{r}(S^{*}),x,y)]+2C_{3}+(1-\beta)\cdot C_{1}+(1-\gamma)\cdot C_{2}$ (34)
$\leq\mathbb{E}_{(x,y)\sim\mathcal{T}}[L(M_{r}(S^{*}),x,y)]+2C_{3}+(1-\beta\gamma)\cdot(C_{1}+C_{2}),$ (35)
where:
  • $C_{2}=\sum_{k=1}^{K}w_{k}\cdot\Delta_{k}$ represents the maximum possible improvement across all clusters,
  • $C_{3}=\sum_{k=1}^{K}w_{k}\sum_{j\neq k}\delta_{kj}$ represents the total cross-cluster effects.
Step 7: Let $C=C_{1}+C_{2}$, considering $\delta_{kj}\to 0$ as claimed in the Assumption 2,
$\mathbb{E}_{(x,y)\sim\mathcal{T}}[L(M_{r}(S_{PASER}),x,y)]\leq\mathbb{E}_{(x,y)\sim\mathcal{T}}[L(M_{r}(S^{*}),x,y)]+(1-\beta\gamma)\cdot C.$ (36)
C.4 Interpretation of the Error Bound
The error bound provides several key insights, explicitly highlighting the role of each assumption:
Performance Decomposition (Assumption 1): This foundational assumption enables modular analysis by decomposing model performance into distinct capabilities, allowing PASER to target specific degraded capabilities with focused recovery resources.
Capability Independence (Assumption 2): By establishing bounded cross-capability influence, this assumption validates PASER’s independent optimization within each capability cluster, ensuring that improvements in one area don’t significantly interfere with others.
Conceptual Consistency Monotonicity (Assumption 3): By ensuring monotonicity within conceptually consistent subsets, PASER’s CCG mechanism effectively mitigates negative tuning effects, making the selection problem more tractable.
Sample Selection Submodularity (Assumption 4): The approximated submodularity enables efficient greedy-like selection methods to achieve near-optimal results within each capability cluster.
Intra-Capability Selection Effectiveness (Assumption 5): The parameter $\gamma$ reflects PASER’s efficiency in selecting the most valuable samples within each capability cluster. When $\gamma$ approaches 1, PASER’s intra-cluster selection approaches optimality.
Inter-Capability Budget Allocation Effectiveness (Assumption 6): The parameter $\beta$ reflects PASER’s efficiency in allocating the budget across different capability clusters. When $\beta$ approaches 1, PASER’s budget allocation approaches the theoretical optimal allocation.
Combined Effect: The error term $(1-\beta\gamma)$ indicates that when both selection effectiveness and allocation effectiveness are high, PASER’s performance approaches the global optimal solution.
This theoretical framework explains PASER’s strong empirical performance: by intelligently allocating budget based on capability degradation degree (improving $\beta$) and selecting the most valuable samples within each cluster while maintaining concept consistency (improving $\gamma$), PASER achieves near-optimal recovery performance.
@@ C.4 Interpretation of the Error Bound
Appendix D Evaluation on Mathematical Problem Solving Tasks
  • GSM8K [Cobbe et al., 2021]: A dataset containing 8.5K high-quality grade school math word problems that test various mathematical problem solving capabilities, including arithmetic, algebra, and word problem solving.
  • Minerva Math [Lewkowycz et al., 2022]: A comprehensive mathematical evaluation dataset covering diverse topics in mathematics ranging from arithmetic to calculus, with problems requiring multi-step reasoning.
@@ Appendix D Evaluation on Mathematical Problem Solving Tasks
Appendix E Evaluation on Code Generation Tasks
@@
  • HumanEval [Chen et al., 2021]: A widely-used code generation benchmark consisting of 164 hand-crafted Python programming problems that test various programming concepts. Each problem contains a function signature with docstring and test cases, requiring models to complete the implementation. The benchmark evaluates functional correctness using the Pass@$k$ metric, which measures the percentage of problems where a correct solution appears in k samples.
  • MBPP [Austin et al., 2021]: A programming benchmark containing 974 Python programming problems focused on basic programming tasks. Each problem includes a natural language description, test cases, and a reference solution, making it particularly suitable for evaluating language-to-code generation capabilities. MBPP uses the same Pass@$k$ evaluation metric as HumanEval but generally features simpler problems with a broader coverage of basic programming concepts.
@@
Appendix F Extended Experiments on Post-quantization Recovery Training
Though the method descriptions and the experiments in the main body are mainly around the LLM pruning scenario, our PASER framework can actually be extended seamlessly to other LLM compression scenario. To further demonstrate its applicability, we conduct the experiments on post-quantization recovery training and compare our PASER with corresponding instruction tuning data selection baselines. In detail, we choose two most representative methods: Round-To-Nearest (RTN) [Frantar and Alistarh, 2022, Yao et al., 2022], GPTQ [Frantar et al., 2023], to perform the LLM quantization. It should be clarified that RTN method, which rounds all weights to the nearest quantized value on exactly the same asymmetric per-row grid, is actually the fundamental technique in most works about LLM quantization [Frantar and Alistarh, 2022, Yao et al., 2022, Park et al., ]. Its runtimes scales well to the models with many billion parameters due to the direct rounding. According to the results provided in Table 7 and 8, we can observe that the PASER can still effectively enhance the recovery performance and outperform the data selection baselines on averaged reasoning performance and zero-shot perplexity for both LLaMA2-7B and LLaMA2-13B models. Meanwhile, recovery data selection baselines can indeed achieve the stronger performance than full data and random baselines, which validates the necessity of conducting recovery data selection even in the LLM quantization scenario. Furthermore, comparing these results with Table 1 and 2, it can be noticed that the improvement space of PASER in Table 7 and 8 has been reduced to some extent. This is because the post-compression performance of such quantization schemes has been competitive enough, which can reflected from the w/o training row.
@@ Appendix F Extended Experiments on Post-quantization Recovery Training
Appendix G Extended Experiments on Recovery Training with Knowledge Distillation
Inspired by  [Muralidharan et al., 2024], we explore the knowledge distillation as the recovery post-training paradigm instead of the standard supervised learning with the groundtruth label. Here, we set the original model $M_{o}$ as the teacher and the pruned model $M_{p}$ as the student. The mean KL divergence [Kullback and Leibler, 1951] between the output probability distribution of $M_{o}$ and that of $M_{p}$ is taken as the objective function. Comparing the corresponding results under different pruning schemes in Table 9 with that in Table 1, we can first observe that knowledge distillation can effectively improve the recovery performance on both reasoning and language modeling tasks in most cases. In particular, the reasoning performance of PASER is improved by 0.348 points on average among such four pruning schemes. Interestingly, the knowledge distillation recovery performance of Full Data under LLM-Pruner is much better than that with standard label-supervised learning. This demonstrates that knowledge distillation is also a promising approach to avoid the misleading information from the irrelevant or conflicting samples existing in the original dataset. Because its learning process directly imitates the unpruned model behavior instead of the provided labels, thus better preserving the thinking and decision-making consistency with the original model. As a summary, distilling the knowledge of unpruned model into the pruned model can be regarded as an effective way to enhance the recovery performance, though bring more memory overhead. Furthermore, stronger layer-wise distillation can also be taken into consideration [Jiao et al., 2020].